A private university charges the same tuition for both in-state and out-of-state students, and it notices that in-state and out-of-state students seem to respond differently to tuition changes. The quantity demanded for in-state and out-of-state students at different tuition levels is provided below Tuition Quantity deded Quantity demanded nin-state applicants) (out-of-state applicants) $10,000 15,000 $20,000 $30,000 6,000 5,000 4,000 3,000 12,000 9,000 6,000 3,000 As the price of tuition rises from $10000 to $15000, the price elasticity of demand for tuition for out-of-state applicants is than it is for in-state applicants elastic Choose one: o A. less
$15,000 $20,000 $30,000 5,000 4,000 3,000 9,000 6,000 3,000 As the price of tuition rises from $10000 to $15000, the price elasticity of demand for tuition for out-of-state applicants is than it is for in-state applicants. elastic Choose one: A. less o B. more Part 2 (2 points) What is the in-state elasticity and the out-of-state elasticity? Give your answer to two decimals. In state: Out of state:
Answer #1

answer:

Part 1- As the price of tuition rises from $10,000 to $15,000, the price elasticity of demand for tuition for out - of - state applicants is more elastic than it is for in - state applicants  

Part 2- the in-state elasticity is :0.33

The out-of-state elasticity is : 0.50

Explanation:

formula of price elasticity of demand = % change in quantity demanded / % change in price

% change in quantity demanded = (change in quantity/ initial quantity) *100

% change in price = change in price/ initial price *100

price elasticity of demand for tuition for in state applicants is :

price has increased from $10000 to $15000

change in price = $5000

initial price = $10000

% change in price = 5000/ 10000 *100 = 50%

quantity demanded has decreased from 6000 to 5000

change in quantity demanded = 1000

initial quantity demanded = 6000

% change in quantity demanded = (6000 - 5000/ 6000) *100 = 16.67 %

price elasticity of demand = 16.67 / 50 = 0.33

so price elasticity of demand for in-state students is 0.33

price elasticity of demand for out -of- state students :

price increased from $10000 to $15000

quantity demanded decreased from 12000 to 9000

% change in quantity demanded = 12000 -9000 / 12000 *100

% change in quantity demanded = 3000/12000 * 100 = 25%

% change in price = 50%

price elasticity of demand = 25 / 50 = 0.5

so price elasticity of demand for out - of -state students is 0.5